1/1

Lord George Gordon Byron ,

She Walks in Beauty

She walks in Beauty, like the Night Of cloudless climes and starry skies; And all that’s best of dark and bright Meet in her aspect and her eyes; Thus mellowed to the tender light Which Heaven to gaudy day denies –

One shade the more – one ray the less Had half impaired the nameless grace Which waves in every raven tress Or softly lightens o’er her face Where thoughts serenely sweet express How pure – how dear their dwelling place!

And on that cheek, and o’er that brow, So soft – so calm – yet eloquent, The smiles that win, the tints that glow, But tell of days in goodness spent – A mind at peace with all below – A Heart – whose love is innocent!

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1109-2238 (2015-01-17 17:34:42)

Tiu ĉi poemo troviĝas en http: // petercochran. files. wordpress. com/ 2009/ 03/ hebrew_melodies. pdf .

http://www.poezio.net/version?poem-id=1109&version-id=,2238